Maximize St. 65x + 45Y 5x + 4 Y J 40 Ixt OY L 15 Ixt I Y L 22 X Y Z
Answer #1

22 18 16 12 2 14 16 18 20 1o 12

The objective function is

Max [ 65x+45Y]

constraints

5X+4Y=>40,

x<=15

x+y<=22

x=>0, y=>0

Feasible region is shown by the graph above, with points of optimality ( 0,22) (0.10) ( 15.7) and (22,0)

Value of objective fnction at

(0,22) = 65x0+45x22 = 990

(0,10) =10x45 =450

(15,7) = 15x65+7x45 =1290

(8,0) = 8x65 = 520

The optimal solution is x=15, y =7 with maximum value of 1290.

Note: I have answered same problem before. If it is the same person posting it again, please let me know if something needs to be explained.